Factorise each of the following:
a 2m − 8
b xy − xw
c −3g − 27
d 12x + 8 − 16a
e 4w2 − 12w
f 5k + hk2
1
Expert's answer
2011-05-26T06:37:49-0400
a. 2m - 8 = 2(m-4)
b. xy - xw = x(y - w)
c. -3g - 27 = -3 ( g + 7)
d. 12x + 8 - 16a = 4(3x + 2 - 4a)
e. 4w2 - 12w = 4w ( w - 3)
f. 5k + hk2 = k ( 5 + hk)